In triangle ABC, Anle A is larger than angle C and smaller then angle B by the same amount.If angle B is 67,angle c is?
Answer by fcabanski(1391)

Anle A is larger than angle C and smaller then angle B by the same amount means:
A-C=B-A and since B=67 that becomes A-C=67-A.


Also known is that A+67+C=180. Manipulate this to get A in terms of C, so we can substitute that back into the first equation and then solve it for C.


A=180-67-C


180-67-C-C=67-180+67+C
113-2C=-46+C
-3C=-159
C=53

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com